OpenStudy (anonymous):

Why is the exterior angle of a triangle equal to the sum of the two remote interior angles?

OpenStudy (anonymous):



OpenStudy (anonymous):

I can't see the whole drawing,,, @agostino

OpenStudy (anonymous):



OpenStudy (anonymous):

In words, consider a triangle with angles a, b, c. Say we extend a side of the triangle such that there is another angle d which satisfies d=180-c, i.e. an exterior angle of the triangle. We have a+b+c=180 and by c+d=180, we have a+b=d.

OpenStudy (anonymous):

\[A+B+C=180\]PROPERTY OF TRIANGLE SUM\[D=180-C\]SUPPLEMENTARY ANGLES\[NOW C=180-D \]

OpenStudy (anonymous):

SUBSTITUTE IT INTO THE 1ST EQUATION

OpenStudy (anonymous):

A+B+(180-D)=180 A+B+180-D=180 A+B=D

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

HENCE PROVED!!